OpenStudy (newwar):

The function f(x)=5000(0.98)0.3x represents the number of white-blood cells, per cubic millimeter, in a patient x days after beginning treatment for a virus. What is the average decrease per day in white-blood cells per cubic millimeter between days 1 and 5? ​ −29.75 mm3/day ​ ​ −34.75 mm3/day ​ −282.47 mm3/day −353.08 mm3/day

OpenStudy (3mar):

Do you think any idea?

OpenStudy (newwar):

i think it is C

OpenStudy (3mar):

May I see your work guided you to that answer, please?

OpenStudy (newwar):

to mine

OpenStudy (3mar):

\[\Large f(x)=5000(0.98)^{0.3x}\] That is the main equation of the decaying of white-blood cells number. the average decay could be calculated as follows: \[\frac{\left(f\left(5\right)-f\left(1\right)\right)}{5-1}\] https://www.desmos.com/calculator/fxlmllovm8